OpenStudy (anonymous):

find critical numbers and point of inflection f(x)=x^2-2x +1 /X+1

OpenStudy (anonymous):

\[f \prime=2x-2-1/(x+1) \] crit numbers \[\sqrt{3/2}\] and\[-\sqrt{3/2}\] inflection pt x=1/2